# Uncertainty Quantification for Prior-Data Fitted Networks using Martingale Posteriors

### v2

@@ 1 INTRODUCTION
Prior-data fitted networks (PFNs) are foundation models (Hollmann et al.,, 2023; Müller et al.,, 2022) that allow for in-context learning, i.e., the ability to learn at inference time without any parameter updates (Garg et al.,, 2022). TabPFN, a transformer pre-trained on synthetic data for in-context learning on tabular datasets, has recently attracted a lot of interest. TabPFN (Hollmann et al.,, 2023, 2025) and extensions such as TuneTables (Feuer et al.,, 2024) and LocalPFN (Thomas et al.,, 2024) have been shown to achieve state-of-the-art performance on tabular benchmarks by pre-training on purely synthetic data. Since PFNs and extensions learn in-context, there is no need for further model (fine-)tuning on the inference task.
Recent extensions of PFNs allow their applicability to large datasets (Feuer et al.,, 2024), the use of PFN “priors” for latent variable models (Reuter et al.,, 2025), and simultaneously minimizing bias and variance to improve their performance (Liu and Ye,, 2025). PFNs are also related to simulation-based inference and amortized inference but have slightly different goals and do not amortize across a single but multiple datasets (Reuter et al.,, 2025). While introduced as a Bayesian method and approximation to the posterior predictive, PFNs can also be interpreted as pre-tuned untrained predictors (Nagler,, 2023). This also relates to the question of what uncertainty PFN models can provide.
@@
We propose a principled and efficient method to construct Bayesian posteriors for such estimates using the idea of Martingale Posteriors (MPs). In particular:
@@
    We introduce a new extension of the MP framework of Fong et al., (2023) for inference of predictive quantities conditional on a specific feature value $x$.
@@
Our work provides an essential tool for principled inference with the exponentially popular PFN methods. A brief discussion of potential extensions to large language models is provided in Section 5.
Refer to caption Refer to caption Refer to caption
@@ 2.1 Prior-Data Fitted Networks
Prior-data fitted networks are foundation models trained to approximate the posterior predictive density
$\operatorname{PPD}(y|x)=p(y|x,\mathcal{D}_{n}),$
which quantifies the likelihood of observing label $y$ given that the feature is $x$ and $\mathcal{D}_{n}$ has been observed. The PPD is a Bayesian concept and implicitly involves a prior over the distributions $P$ that could have generated the data. To approximate the PPD with a PFN, a deep neural network—typically a transformer—is pre-trained on simulated datasets with diverse characteristics. After pre-training, the network weights are fixed, and the approximate PPD for a new training set can be computed through a single forward pass without additional training or tuning. The PPD quantifies uncertainty about the label $y$. However, it mixes the aleatoric and epistemic components of uncertainty (Hüllermeier and Waegeman,, 2021). From the PPD alone, it is impossible to disentangle these parts. Consequently, PFNs do not provide uncertainty estimates for predictive summaries, such as conditional mean, probabilities, and quantiles (see Figure 1).
@@ 2.2 Bayesian Inference
$\Pi(\mu(x)\in A)=\int\mathds{1}\left\{\int y\,p_{\theta}(y|x)dy\in A\right\}\pi(\theta|\mathcal{D}_{n})\,\,d\theta,$
@@ 2.3 Martingale Posteriors
Martingale posteriors were recently introduced by Fong et al., (2023) as a new method for Bayesian uncertainty quantification. Its core idea is to reverse the direction of the Bayesian inference. In classical Bayesian inference, the posterior is derived from a prior and likelihood, which then implicitly leads to the PPD. MP inference starts from the PPD and leaves the prior $\pi(\theta)$ implicit. An appropriate sampling scheme and Doob’s theorem then allow us to derive posteriors for virtually all quantities of interest (e.g., the conditional mean $\mu(x)$).
@@
$F_{\infty}(y)=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(y_{n+i}\leq y),$
@@
$\Pi(\theta\in A|\mathcal{D}_{n})=\int\mathds{1}\{\theta(F_{\infty})\}d\Pi(F_{\infty}|\mathcal{D}_{n}),$
where $A$ is any Borel set on the space where the parameter $\theta$ lives. Furthermore, Doob’s theorem (Doob,, 1949) implies that $\Pi(\theta|\mathcal{D}_{n})$ coincides with the classical Bayes posterior for the prior $\pi(\theta)$ implicit in the PPD (Fong et al.,, 2023).
@@ 3 EFFICIENT MARTINGALE POSTERIORS FOR PRIOR-DATA FITTED NETWORKS
Martingale posteriors allow for Bayesian inference directly from the PPD. PFNs approximate the PPD, so using PFNs to construct a martingale posterior seems natural. However, there are two problems. First, modern PFNs are based on transformer architectures that require $\Omega(n^{2})$ operations for a forward pass on a training set size of $n$. Iteratively computing $p(y|y_{1:(n+k)})$ for $k=1,\dots,N$ thus has complexity $\Omega(N^{3})$, which is prohibitive. Second, Falck et al., (2024) found that modern transformer-based LLMs substantially deviate from the martingale property
@@
LABEL:{fig:martprop} (experiment description in Section 4.1) shows a QQ-plot of the average CDF resulting from PFN’s PPD at iteration 0 vs. the CDF after $k$ forward iterations of the sampling procedure. The martingale property is satisfied when this distribution remains constant, i.e., all lines lie on the $x=y$ diagonal. However, we see an apparent change in the distribution over the different sampling horizons. This reinforces that PFNs approximate Bayesian inference only in a weak sense, even though they were conceived from Bayesian ideas (Hollmann et al.,, 2023; Müller et al.,, 2021).
Without the martingale property, the MP sampling procedure leads to meaningless results. Instead, we propose to use the PPD implied by the PFN only as a starting point for the sampling scheme. This PPD is then iteratively updated using the nonparametric Gaussian copula approach of Fong et al., (2023), which ensures the martingale property. The resulting martingale posterior effectively treats the PFN output as an informed prior, without requiring the PFN to provide coherent posterior updates.
@@ 3.1 Martingale Posteriors for Conditional Inference
We extend the unconditional sampling scheme outlined in the previous section to the conditional inference setting. Fong et al., (2023) already proposed one such extension. Their scheme involves forward sampling of the features $x_{(n+1):(n+N)}$. The distribution of the features isn’t of primary interest but significantly complicates the sampling procedure, which the authors resolved through heuristic simplifications. In contrast to Fong et al., (2023), we propose to sample only the labels $y_{(n+1):(n+N)}$ conditional on the event that $x_{n+k}=x$, for a fixed value of $x$ and all $k=1,\dots,N$.
@@
Denote the long-run empirical distribution of the obtained sample by
$F_{\infty,x}=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(y_{n+i}\leq y),$
which is again a random function, even in the limit. Repeating the iterative sampling procedure gives us its distribution $\Pi(F_{\infty,x}|\mathcal{D}_{n})$. For any conditional parameter $\theta(x)=\theta(P(\cdot\,|x))$ of interest, the martingale posterior is now given as
$\Pi(\theta(x)\in A|\mathcal{D}_{n})=\int\mathds{1}\{\theta(F_{\infty,x})\}d\Pi(F_{\infty,x}|\mathcal{D}_{n}).$
Common examples of the parameter $\theta(x)$ are the conditional mean $\theta(x)=\int y\,dP(y|x)dy$ or a conditional $\varrho$-quantile $\theta(x)=P^{-1}(\varrho|x)$.
@@ 3.2 Efficient PPD Updates Based on the Gaussian Copula
Observe that $p_{0}(y)=p(y|x,\mathcal{D}_{n})$ is the PPD approximated by the PFN $g_{\theta}$. However, the following update distributions $p_{1},p_{2},\dots$ are generally intractable. To alleviate this, Fong et al., (2023) proposed a computationally efficient, nonparametric method based on Dirichlet Process Mixture Models (DPMMs) and a copula decomposition of the conditional $p_{k}$. Specifically, we set
@@
and $\Phi$ is the standard normal cumulative distribution function. The method has a hyperparameter $\rho$, corresponding to a bandwidth that smoothes the updates. Fong et al., (2023) proposed to tune this by maximizing the likelihood of the updated densities over the observed data. To increase the alignment of the updates with the PFN baseline, we simulate a new sample from the PFN and optimize the bandwidth on the simulated data. More details can be found in Section B.1
@@ 3.3 Flexible Learning Rate Schedules
Fong et al., (2023) suggested the learning rate $\alpha_{i}=(2-i^{-1})(i+1)^{-1}$ for the unconditional context, where distribution functions can be estimated with convergence rate $n^{-1/2}$. As we shall see in Proposition 3 below, the learning rate $\alpha_{i}$ determines the contraction rate of the posterior. Our setting, however, is conditional on $x\in\mathbb{R}^{d}$, where nonparametric methods such as TabPFN must converge slower than $n^{-1/2}$, with rates depending on the dimension and smoothness of the estimation problem (see, e.g., Stone,, 1982).
@@
where $C$ and $\beta$ should be chosen such that the behavior of the update (3) roughly matches the behavior of the PFN. In our experiments, we implement the following three schedules
  • default: $C=2$, $\beta=1$ as in (Fong et al.,, 2023),
@@
The constants $C$ are set to normalize the schedules to a comparable starting scales. We refer to our discussion after Proposition 3 in the next section for a motivation of the exponent $\beta$, and to Section 4.4 for empirical evaluation.
@@ 3.4 Theoretical Properties
Despite the simplicity of the updates given in (3), they provide essential theoretical guarantees. The following is a direct consequence of Theorem 3 by Fong et al., (2023).
Proposition 1.
@@
By de Finetti’s theorem (e.g., Theorem 1.49 in Schervish,, 2012), it then follows that there is a random variable $\Theta$ such that $(z_{1},z_{2},\dots)$ is conditionally iid given $\Theta$. The distribution of $\Theta$ can be interpreted as an implicit prior. In our setting, this prior depends both on the initial PPD implied by the PFN and the Gaussian copula updates specified by (3). In particular, the following result follows from Proposition 1 above and Theorem 2.2 of Berti et al., (2004).
Proposition 2.
@@
$|P_{\infty,x}(y)-P_{N}(y)|\leq C\sqrt{\frac{\log(2/\delta)}{2\beta-1}}(n+N)^{-\beta+1/2}.$
The proof, given in Appendix A, is based on a time-uniform version of the Azuma-Hoeffding concentration inequality (Howard et al.,, 2020), the martingale property of the copula updates, and the learning rate schedule. Importantly, time-uniform martingale concentration allows us to directly bound the deviation of interest $|P_{\infty,x}-P_{N}|$ as opposed to $|P_{M}-P_{N}|$ with $M\geq N$ as in Fong et al., (2023).
Proposition 3 quantifies how well the distribution $P_{N}$ approximates $P_{\infty,x}$ after $N$ updates. Setting $N=0$ corresponds to the case where $P_{N}$ equals the initial PPD given only the observed data. In our setting, this is the output from the PFN. The random limit $P_{\infty,x}$ fluctuates around this point estimate, and the spread of these fluctuations characterizes the spread of the martingale posterior for $\theta(P(\cdot|x))$.
@@ 3.5 The Role of $\beta$
The case $\beta=1$ corresponds to the choice $\alpha_{i}\sim i^{-1}$ proposed by Fong et al., (2023) in the unconditional context with contraction rate of $n^{-1/2}$. Contraction rates for nonparametric regression problems are typically much slower than $n^{-1/2}$ (Ghosal and Van der Vaart,, 2017, Chapter 9), however. To accommodate such slow contraction rates, it is therefore crucial to allow for learning rates with $\beta<1$. The choice $\beta=d/(d+4)$ of our type 1 schedule corresponds to the optimal contraction rate we can expect in a nonparametric regression problem with $d$ features and twice continuously differentiable regression curve. However, Nagler, (2023) demonstrated that current PFN architectures yield inconsistent estimators, as their bias does not vanish as $n\to\infty$. This suggests that to get valid inferences, the posterior should never contract fully. This happens in the regime $\beta\in(0,1/2]$, where the bound in Proposition 3 becomes vacuous, but the algorithm still converges to a random limit $P_{\infty,x}$ because of Proposition 2. This motivates the choice $\beta=4/(d+4)$, which has $\beta\leq 1/2$ for the regime $d\geq 4$, where the bias becomes notable already on moderate sample sizes.

### v3

@@ 1 Introduction
Prior-data fitted networks (PFNs) are foundation models (Hollmann et al., 2023; Müller et al., 2022) that allow for in-context learning, i.e., the ability to learn at inference time without any parameter updates (Garg et al., 2022). TabPFN, a transformer pre-trained on synthetic data for in-context learning on tabular datasets, has recently attracted a lot of interest. TabPFN (Hollmann et al., 2023, 2025) and extensions such as TuneTables (Feuer et al., 2024) and LocalPFN (Thomas et al., 2024) have been shown to achieve state-of-the-art performance on tabular benchmarks by pre-training on purely synthetic data. Since PFNs and extensions learn in-context, there is no need for further model (fine-)tuning on the inference task.
Recent extensions of PFNs allow their applicability to large datasets (Feuer et al., 2024), the use of PFN “priors” for latent variable models (Reuter et al., 2025), and simultaneously minimizing bias and variance to improve their performance (Liu and Ye, 2025). PFNs are also related to simulation-based inference and amortized inference but have slightly different goals and do not amortize across a single but multiple datasets (Reuter et al., 2025). While introduced as a Bayesian method and approximation to the posterior predictive, PFNs can also be interpreted as pre-tuned untrained predictors (Nagler, 2023). This also relates to the question of what uncertainty PFN models can provide.
@@
We propose a principled and efficient method to construct Bayesian posteriors for such estimates using the idea of Martingale Posteriors (MPs) proposed in Fong et al. (2023). In particular:
@@
    We introduce a new extension of the MP framework for inference of predictive quantities conditional on a specific feature value $x$.
@@
  5. 5.
    We perform a variety of ablation studies providing insights into the driving factors of our proposal’s efficacy and failure modes of alternative methods.
Our work provides an essential tool for principled inference with the exponentially popular PFN methods and overcomes limitations of previous methods that struggle with complex datasets or PFN’s missing martingale property (cf. Figure 1). A brief discussion of potential extensions to large language models is provided in Section 5.
Figure 1: Comparison of estimated median function (orange line) and credible intervals (shaded blue area) for a non-linear feature effect (dashed blue line) in a Bayesian additive model based on different methods (facets). Numbers in brackets after each method state the marginal miscoverage, i.e., the absolute difference between the method’s average coverage and the nominal level of 90% (smaller is better). Conditional coverages for individual $x_{1}$ values are provided in the inset plots with the nominal level marked as a gray dashed line. “Optimal” corresponds to the posterior induced by the chosen prior and likelihood, which in this case can be analytically derived (see Section 4.3 for details). Its miscoverage is 0.018 (which is plausible, being in a small data setting) and very close to our method.
@@ 2.1 Prior-Data Fitted Networks
Prior-data fitted networks are foundation models trained to approximate the posterior predictive density $\operatorname{PPD}(y|x)=p(y|x,\mathcal{D}_{n})$, which quantifies the likelihood of observing label $y$ given that the feature is $x$ and $\mathcal{D}_{n}$ has been observed. The PPD is a Bayesian concept and implicitly involves a prior over the distributions $P$ that could have generated the data. To approximate the PPD with a PFN, a deep neural network—typically a transformer—is pre-trained on simulated datasets with diverse characteristics. After pre-training, the network weights are fixed, and the approximate PPD for a new training set can be computed through a single forward pass without additional training or tuning.
The PPD quantifies uncertainty about the label $y$. However, it mixes the aleatoric and epistemic components of uncertainty (Hüllermeier and Waegeman, 2021). From the PPD alone, it is impossible to disentangle these parts. Consequently, PFNs do not provide uncertainty estimates for predictive summaries, such as conditional mean, probabilities, and quantiles.
@@ 2.2 Bayesian Inference
$\Pi(\mu(x)\in A)=\int\bm{1}\left\{\int y\,p_{\theta}(y|x)dy\in A\right\}\pi(\theta|\mathcal{D}_{n})\,\,d\theta,$
@@ 2.3 Martingale Posteriors
Martingale posteriors were recently introduced by Fong et al. (2023) as a new method for Bayesian uncertainty quantification. Its core idea is to reverse the direction of the Bayesian inference. In classical Bayesian inference, the posterior is derived from a prior and likelihood, which then implicitly leads to the PPD. MP inference starts from the PPD and leaves the prior $\pi(\theta)$ implicit. An appropriate sampling scheme and Doob’s theorem then allow us to derive posteriors for virtually all quantities of interest (e.g., the conditional mean $\mu(x)$).
@@
$F_{\infty}(y)=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\bm{1}(y_{n+i}\leq y),$
@@
$\Pi(\theta\in A|\mathcal{D}_{n})=\int\bm{1}\{\theta(F_{\infty})\}d\Pi(F_{\infty}|\mathcal{D}_{n}),$
where $A$ is any Borel set on the space where the parameter $\theta$ lives. Furthermore, Doob’s theorem (Doob, 1949) implies that $\Pi(\theta|\mathcal{D}_{n})$ coincides with the classical Bayes posterior for the prior $\pi(\theta)$ implicit in the PPD (Fong et al., 2023).
@@ 3 Efficient Martingale Posteriors for Prior-Data Fitted Networks
Martingale posteriors allow for Bayesian inference directly from the PPD. PFNs approximate the PPD, so using PFNs to construct a martingale posterior seems natural. However, there are two problems. First, modern PFNs are based on transformer architectures that require $\Omega(n^{2})$ operations for a forward pass on a training set size of $n$. Iteratively computing $p(y|y_{1:(n+k)})$ for $k=1,\dots,N$ thus has complexity $\Omega(N^{3})$, which is prohibitive. Second, Falck et al. (2024) found that modern transformer-based LLMs substantially deviate from the martingale property
@@
LABEL:{fig:martprop} (experiment description in Section 4.1) shows a QQ-plot of the average CDF resulting from PFN’s PPD at iteration 0 vs. the CDF after $k$ forward iterations of the sampling procedure. The martingale property is satisfied when this distribution remains constant, i.e., all lines lie on the $x=y$ diagonal. However, we see an apparent change in the distribution over the different sampling horizons. This reinforces that PFNs approximate Bayesian inference only in a weak sense, even though they were conceived from Bayesian ideas (Hollmann et al., 2023; Müller et al., 2021).
Without the martingale property, the MP sampling procedure leads to meaningless results. Instead, we propose to use the PPD implied by the PFN only as a starting point for the sampling scheme. This PPD is then iteratively updated using the nonparametric Gaussian copula approach of Fong et al. (2023), which ensures the martingale property. The resulting martingale posterior effectively treats the PFN output as an informed prior, without requiring the PFN to provide coherent posterior updates.
@@ 3.1 Martingale Posteriors for Conditional Inference
We extend the unconditional sampling scheme outlined in the previous section to the conditional inference setting. Fong et al. (2023) already proposed one such extension. Their scheme involves forward sampling of the features $x_{(n+1):(n+N)}$. The distribution of the features isn’t of primary interest but significantly complicates the sampling procedure, which the authors resolved through heuristic simplifications. In contrast to Fong et al. (2023), we propose to sample only the labels $y_{(n+1):(n+N)}$ conditional on the event that $x_{n+k}=x$, for a fixed value of $x$ and all $k=1,\dots,N$.
@@
Denote the long-run empirical distribution of the obtained sample by $F_{\infty,x}=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\bm{1}(y_{n+i}\leq y)$, which is again a random function, even in the limit. Repeating the iterative sampling procedure gives us its distribution $\Pi(F_{\infty,x}|\mathcal{D}_{n})$. For any conditional parameter $\theta(x)=\theta(P(\cdot\,|x))$ of interest, the martingale posterior is now given as
$\Pi(\theta(x)\in A|\mathcal{D}_{n})=\int\bm{1}\{\theta(F_{\infty,x})\}d\Pi(F_{\infty,x}|\mathcal{D}_{n}).$
Common examples of the parameter $\theta(x)$ are the conditional mean $\theta(x)=\int y\,dP(y|x)dy$ or a conditional $\varphi$-quantile $\theta(x)=P^{-1}(\varphi|x)$.
@@ 3.2 Efficient PPD Updates Based on the Gaussian Copula
Observe that $p_{0}(y)=p(y|x,\mathcal{D}_{n})$ is the PPD approximated by the PFN $g_{\theta}$. However, the following update distributions $p_{1},p_{2},\dots$ are generally intractable. To alleviate this, Fong et al. (2023) proposed a computationally efficient, nonparametric method based on Dirichlet Process Mixture Models (DPMMs) and a copula decomposition of the conditional $p_{k}$. Specifically, we set
@@
and $\Phi$ is the standard normal cumulative distribution function. The method has a hyperparameter $\rho$, corresponding to a bandwidth that smoothes the updates. Fong et al. (2023) proposed to tune this by maximizing the likelihood of the updated densities over the observed data. To increase the alignment of the updates with the PFN baseline, we simulate a new sample from the PFN and optimize the bandwidth on the simulated data. More details can be found in Section B.1
@@ 3.3 Flexible Learning Rate Schedules
Fong et al. (2023) suggested the learning rate $\alpha_{i}=(2-i^{-1})(i+1)^{-1}$ for the unconditional context, where distribution functions can be estimated with convergence rate $n^{-1/2}$. As we shall see in Proposition 3.3 below, the learning rate $\alpha_{i}$ determines the contraction rate of the posterior. Our setting, however, is conditional on $x\in\mathbb{R}^{d}$, where nonparametric methods such as TabPFN must converge slower than $n^{-1/2}$, with rates depending on the dimension and smoothness of the estimation problem (see, e.g., Stone, 1982).
@@
where $C$ and $\beta$ should be chosen such that the behavior of the update (3) roughly matches the behavior of the PFN. In our experiments, we implement the following schedules:
  • default: $C=2$, $\beta=1$ as in (Fong et al., 2023),
@@
The constants $C$ are set to normalize the schedules to a comparable starting scales. We refer to our discussion after Proposition 3.3 in the next section for a motivation of the exponent $\beta$, and to Section 4.4 for empirical evaluation.
@@ 3.4 Theoretical Properties
Despite the simplicity of the updates given in (3), they provide essential theoretical guarantees. The following is a direct consequence of Theorem 3 by Fong et al. (2023).
Proposition 3.1.
@@
By de Finetti’s theorem (e.g., Theorem 1.49 in Schervish, 2012), it then follows that there is a random variable $\Theta$ such that $(z_{1},z_{2},\dots)$ is conditionally iid given $\Theta$. The distribution of $\Theta$ can be interpreted as an implicit prior. In our setting, this prior depends both on the initial PPD implied by the PFN and the Gaussian copula updates specified by (3). In particular, the following result follows from Proposition 3.1 above and Theorem 2.2 of Berti et al. (2004).
Proposition 3.2.
@@
$|P_{\infty,x}(y)-P_{N}(y)|\lesssim\sqrt{\frac{\log(2/\delta)}{2\beta-1}}(n+N)^{-\beta+1/2}.$
The proof, given in Appendix A, is based on a time-uniform version of the Azuma-Hoeffding concentration inequality (Howard et al., 2020), the martingale property of the copula updates, and the learning rate schedule. Importantly, time-uniform martingale concentration allows us to directly bound the deviation of interest $|P_{\infty,x}-P_{N}|$ as opposed to $|P_{M}-P_{N}|$ with $M\geq N$ as in Fong et al. (2023).
Proposition 3.3 quantifies how well the distribution $P_{N}$ approximates $P_{\infty,x}$ after $N$ updates. Setting $N=0$ corresponds to the case where $P_{N}$ equals the initial PPD given only the observed data. In our setting, this is the output from the PFN. The random limit $P_{\infty,x}$ fluctuates around this point estimate, and the spread of these fluctuations characterizes the spread of the martingale posterior for $\theta(P(\cdot|x))$.
3.5 The Role of Initialization
To further understand the role of the initialization in Algorithm 1, the following theoretical statement can be made.
Corollary 3.1.
Let $P_{0,1}$ and $P_{0,2}$ be two PPDs to initialize our algorithm with, and $P_{\infty,x,1}$ and $P_{\infty,x,2}$ be the corresponding martingale limits after infinitely many iterations and $\Delta(y):=P_{\infty,x,1}(y)-P_{\infty,x,2}(y)$. Then for any $x,y$, it holds with probability at least $1-2\delta$ that
$|\Delta(y)|\geq|P_{0,1}(y)-P_{0,2}(y)|-O(n^{-\beta+1/2}).$
This is a direct result of Proposition 3.3 and shows that, up to the contraction rate of the posteriors, the difference in martingale posteriors is as large as the difference in the initializations. Hence, the choice of initialization is the dominating factor for the difference in the martingale posteriors.
@@ 3.6 The Role of $\beta$
The case $\beta=1$ corresponds to the choice $\alpha_{i}\sim i^{-1}$ proposed by Fong et al. (2023) in the unconditional context with contraction rate of $n^{-1/2}$. Contraction rates for nonparametric regression problems are typically much slower than $n^{-1/2}$ (Ghosal and Van der Vaart, 2017, Chapter 9), however. To accommodate such slow contraction rates, it is therefore crucial to allow for learning rates with $\beta<1$. The choice $\beta=d/(d+4)$ of our type 1 schedule corresponds to the optimal contraction rate we can expect in a nonparametric regression problem with $d$ features and twice continuously differentiable regression curve. However, Nagler (2023) demonstrated that current PFN architectures yield inconsistent estimators, as their bias does not vanish as $n\to\infty$. This suggests that to get valid inferences, the posterior should never contract fully. This happens in the regime $\beta\in(0,1/2]$, where the bound in Proposition 3.3 becomes vacuous, but the algorithm still converges to a random limit $P_{\infty,x}$ because of Proposition 3.2. This motivates the choice $\beta=4/(d+4)$, which has $\beta\leq 1/2$ for the regime $d\geq 4$, where the bias becomes notable already for moderate $n$.